# Real-Time Per-Garment Virtual Try-On with Temporal Consistency for Loose-Fitting Garments

### v1

@@ 5. Limitations and Future Work
In addition to the common limitations mentioned by (Wu et al., 2025), such as the necessity of data capture for each garment, the original garment is not removed, and inaccurate composition masks, we would like to highlight two limitations and their future work as follows:
@@
While our method generates visually plausible virtual try-on results for a wide range of body shapes and provides useful styling references (Figure 14), it does not validate the suitability of garment sizing due to the lack of actual user body measurements. This limitation is common among most existing image-based virtual try-on approaches. A promising future direction is to develop a garment synthesis framework that incorporates user-provided body dimensions to enable more accurate and personalized virtual try-on experiences.
@@
Although our recurrent framework incorporates temporal information by conditioning garment synthesis on preceding frames, its primary objective is to enforce inter-frame consistency rather than to model the transient physical dynamics of loose-fitting garments. The datasets we collected using the method of (Wu et al., 2025) comprise sequences of humans rotating in place, offering inadequate motion diversity for capturing the complex physical behaviors of loose garments. To address this limitation, a novel dataset collection method that introduces a broader range of motion while remaining accessible needs to be explored.

### v3

@@ 4.6. Ablation Study on ConvLSTM Module
Figure 16. Ablation study on ConvLSTM module. The difference map visualizing absolute differences between consecutive output frames shows that the ConvLSTM module enhances temporal consistency effectively.
@@ 5. Limitations and Future Work
In addition to the common limitations mentioned by (Wu et al., 2025), such as the original garment not being removed, and inaccurate composition masks, we would like to highlight the following limitations and their future work:
The necessity of per-garment dataset and training.:
Our method requires collecting a dedicated dataset and training a separate network for each garment, which poses scalability challenges. To address this issue, future research could explore a per-garment-type strategy, where a single dataset is collected and a single network is trained for each category of similar garments rather than for each specific item. This approach would substantially reduce training time and improve scalability.
Assumption of tight-fitting garments during inference.:
We assume that users are wearing tight-fitting garments when interacting with our virtual try-on system. The current implementation does not accommodate scenarios where users wear loose-fitting garments, which may lead to performance degradation due to occlusion effects. To overcome this limitation, future research could investigate real-time human body representations that are robust to the presence of loose-fitting garments.
@@
While our method generates visually plausible results for a wide range of body shapes and provides useful styling references (Figure 13), it does not validate the suitability of garment sizing due to the lack of actual user body measurements. Although some general image-based methods (Chen et al., 2023, 2025) incorporate garment size information, they fall short of accurately reflecting the real physical properties of a specific garment. A promising direction for future research is developing a garment synthesis framework that incorporates user-provided body measurements and utilizes per-garment datasets annotated with body measurement metadata. This approach would enable more accurate and personalized virtual try-on experiences.
@@
Although our recurrent framework incorporates temporal information by conditioning garment synthesis on preceding frames, its primary objective is to enforce inter-frame consistency rather than to model the transient physical dynamics of loose-fitting garments, such as fluttering. The datasets we collected using the method of (Wu et al., 2025) comprise sequences of humans rotating in place, offering inadequate motion diversity for capturing the complex physical behaviors of loose garments. To address this limitation, a novel dataset collection method that introduces a broader range of motion while remaining accessible needs to be explored.